OpenStudy (anonymous):

when converting a linear equation into an augmented matrix which equation form is needed

OpenStudy (anonymous):

damn it

OpenStudy (anonymous):

are you on this site all day?

OpenStudy (solomonzelman):

think of a form that has a set up like 4x+3y=0 2x+5x=0

OpenStudy (solomonzelman):

I left my computer on for couple hours, but I wasn't really here.

OpenStudy (solomonzelman):

Maybe not equal to zero, it can be 4x+3y=7 2x+5x=3

OpenStudy (anonymous):

this is gay

OpenStudy (solomonzelman):

standard form.